The​ molarity of glucose solution is 0.02 the mass of glucose in 50cm3(cube) of the solution is what?
1
Expert's answer
2016-04-20T11:00:48-0400
Molarity or molar concentration is defined as the number of moles of solute per liter of solution.
The molar concentration indicated by C(mol/dc3), the amount of substance n(mol), volume V(dc3)
C = n/V
Find the number of moles of glucose:
n(glucose; C6H12O6) = C • V = 0,02 • 50 • 10-3 = 0.001 (mol/dc3)
Then the mass of glucose equals:
m(C6H12O6) = n(C6H12O6) • M(the molar mass of glucose) = 0.001 • 180 = 0.18 (g)
Answer: 0.18 g